# New Coevolution Dynamic as an Optimization Strategy in Group Problem Solving

### v2

@@ Abstract
Coevolution on social models couples the time evolution of the network with the time evolution of the states of the agents. This paper presents a new coevolution dynamic allowing more than one rewiring on the network. We explore how this coevolution can be employed as an optimization strategy for problem-solving capability of task-forces. We used an agent-based model study how this new coevolution dynamic can help a group of agents whose task is to find the global maxima of NK fitness landscapes. Each agent can replace more than one neighbors, and this quantity is a tunable parameter in the model. These rewirings is a way for the agent to obtain information from individuals that were not previously part of its neighborhood. Our results showed that this tunable coevolution can indeed produce gain on the computational cost under certain circunstances. At low average degree (using a random network to connect the agents) and on easy landscape, more rewirings return lower cost, helping the agents to find the global maximum faster. However, at larger average degree and difficult landscape, the effect is more complex. At medium size systems, the coevolution brings optimal results when 3 or 4 neighbors are replaced.
@@ 1 Introduction
In society, it is common to face problems that require collaboration with other people, from everyday challenges to complex tasks, such as group projects at work. In this context, the search for more effective problem-solving strategies becomes a topic of great interest.
Social interactions are inherently complex such that, in recent years, numerous models of social behavior have emerged to explore this intricacy [1, 2, 3] using the framework of Network Science [4]. The nodes of the network represent individuals (or agents) with states defined by the social model while the connections represent the interactions between different individuals. The states of the individuals evolve following the set of rules from the dynamic social model in study. The network itself can also have a time evolution: links can be destroyed or created.
@@
This work is organized as follows: In Section 2.1 we present a brief discussion about the NK model, in section 2.2 we define the computational cost, in section 2.3 we define how the agents are updated, in section 2.4 we define the proposed coevolution mechanism and in section 2.6 we described the random network used to connect the agents. Section 3 is dedicated to present and discuss the results and section 4 is dedicated to our final considerations.
@@ 2.1 NK Fitness Landscapes
The function $f(\mathcal{X})$ has a global maximum $f(\mathcal{X}_{M})=F$ at the string $\mathcal{X}_{M}$. The number of local maxima of the function $f(\mathcal{X})$ is zero at $K=0$ landscape. At $K>0$ the number of local maxima increases. Therefore, at $K=0$, the difference $f(\mathcal{X}_{M})-f(\mathcal{X})$ is proportional to the Hamming distance $d_{H}(\mathcal{X}_{M},\mathcal{X})$ between the vectors [35]. The Hamming distance gives the number of different bits between the two vectors. For example, consider $\mathcal{X}_{1}=(010101)$ and $\mathcal{X}_{2}=(000101)$. The hamming distance is $d_{H}(\mathcal{X}_{1},\mathcal{X}_{2})=1$. At $K>0$, the landscape becomes more complicated because the function $f$ has additional local maxima beside the global one. Thus the difference $f(\mathcal{X}_{M})-f(\mathcal{X})$ is no longer proportional to the hamming distance $d_{H}$. The average number of local maxima increases with $K$ until the limiting case of $K=N-1$, in which the landscape becomes completely random, known as the Random Energy model [36].
@@ 2.2 Computational cost
The efficiency of the search is measured by the time $t_{g}$ in MCS required by the first agent to find the global maximum $F$, in other words, when its string becomes $\mathcal{X}_{M}$ after a number of updates. A small $t_{g}$ means more efficient search: the goal was achieved in less time. So, the lower the better. The normalized computational cost is then defined as [11]:
@@
where $\lambda_{N}$ depends on the value of $N$. $C>C_{I}$ means the interaction between the agents creates more problems than solution, slowing down the quest of the agents for the global maximum. Otherwise, if $C<C_{I}$, the interaction helped the agents making one of them find the global maximum in less time.
@@ 2.3 Strings update
Each agent has its string $\mathcal{X}$ updated when it is selected for analysis. This is how each agent tries to improve its fitness $f(\mathcal{X})$, to eventually reaches the global maximum. We call this agent the target, and he can implement one of the two strategies to perform the update: exploitation or exploration [33]. Exploitation implies the target tries to learn the solution from a neighbour with a higher fitness. This is implemented as a copy, with probability $p$: the target copies one bit of the string of the model agent [35]. Otherwise, exploration means the target aims to innovate a new solution with a better fitness by itself. It is a mutation on the string of the target, and it should happen with probability $1-p$. However if the target has no neighbors or if it is the model of its influence neighborhood, the mutation will happen for sure. If it has no neighbors, it does not have anyone to copy. And if it is already the model, there is no point in copy another agent with a smaller (worst) fitness. The model agent is the one with the highest fitness in the target influence neighborhood. The target agent may be the model if it has a larger fitness than his neighbors.
@@
Both processes change the string of the target agent. The mutation is achieved by a flip of a random bit of the string. The copy process involves more steps: the different bits bewteen the target and the model strings are selected. Then, one of them is randomly selected and flipped on the target string. This is the imperfect copy on the imitation learning context [31]. As a result, the target string gets more similar to the model string (hamming distance decreases). If $K=0$, this means the target agent increases his fitness. On the other hand, if $K>0$, the model one may be a local maxima, and the hamming distance between the target string $\mathcal{X}$ and the global maximum one $\mathcal{X}_{M}$ may increase when the target copies the local maximum (model agent). The target can get trapped on the suboptimal solution (local maximum) [18]. $K=0$ is the easy landscape and increasing $K$ increases the difficulty of the landscape. The most difficult landscape is $K=N-1$ where there is no point on copy because the landscape is completely random.
@@ 2.4 Coevolution mechanism
Figure 1: Schematic illustration of the implemented model. (a) Illustration of one iteration. It starts with the selection of the target agent. $r$ is calculated through Eq. 6. With probability $r^{q}$ the network can suffer a rewiring. The next step is the string update, where a copy happens with probability $p$ and a mutation with probability $1-p$. (b) Illustration of the rewiring process. The target agent (indicated by a dashed black square) loses $L$ neighbors and gets new $L$ ones. The case $L=1$ is illustrated in this picture: the target loses the neighbor on the right and gets acquainted with the agent right above him. The colors of the nodes indicate the components.
The rewiring itself is the temporal evolution of the network. To be considered as coevolution, this network evolution must be coupled to the evolution of the agent string. The constant rewiring (always happening) is not coevolution, although it is a network evolution [35]. As the rewiring itself is a random process, the coupling of the network evolution with the agent states, in our model, will be achieved by the probability of rewiring $r$: it must be calculated considering the fitness of the agents. The rewiring process happens before the string update, as illustrated on Fig. 1(a). Another illustration of one example of the rewiring itself (with $L=1$) is shown Fig. 1(b).
@@ 2.5 Rewiring probability
The goal of each agent is to increase its fitness until one of them, eventually, reaches the global maximum of the landscape. So, having neighbors with better fitness if beneficial. The probability $r$ should be larger when most of the neighbors have smaller fitness, so one of them may be replaced with a higher fitness one. The limit cases of this hypothesis are:
  • the agent is the model: all neighbors have smaller fitness, so it is justifiable to replace some of them. The probability should be maximum: $r=1$.
  • the agent has the smallest fitness of the neighborhood: any copy is already rewarding, so there is no urgency to replace any any neighbor. The probability should be minimum: $r=0$.
  • Otherwise, the probability shoud be such as $10<r<$ 1.
@@
$r=\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}},$ (5)
where $\phi_{t}$ is the fitness of the target agent, $\phi_{s}$ is the smallest fitness and $\phi_{m}$ is the maximum one (fitness of the model) on the influence neighborhood. Notwithstanding, after extensive tests, we did not found any gain on the computational cost with coevolution using this probability.
We have tested several other definitions for the probability, surprinsingly, only one gave interesting results:
$r=1-\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}}.$ (6)
This is the opposite of the first attempt: $r=0$ if the target is the model and $1$ if it has the smallest fitness. The assumption from Eq. 5 is based on the update via copy. It is worth mention that this copy happens with probability $p$, been the other update process the mutation. So, regardless Eq. 5 seems more plausible, it is possible that other mechanisms (such as Eq. 6) may yield better results. After all, the outcome after a mutation is independent if there was a rewiring or not.
For completeness sake, we add one more parameter to the probability, giving it the form $r^{q}$. The free parameter $q\geq 0$ controls the frequency of the rewiring: $q=1$ (probability is just $r$) enables the coevolution and $q=0$ (probability equals $1$) implies a constant rewiring. Possible non-linear effects can be examined using other values subjected to the constraint $q\neq 0$ and $q\neq 1$ [38].
@@ 2.6 Random Network
A network or a graph comprises a collection of vertices interconnected by a set of edges [39, 40]. Each edge, or connection, links a pair of vertices with the number of vertices defining the network size and the set of edges determining its connectivity. Vertices that share an edge are termed neighbors. Throughout this work the terms vertices, agents, and individuals have the same meaning. Various types of networks are used for solutions of physical and biological problems [41, 42, 43]. In this work we will use the Erdös-Rényi (ER) network [44, 45] which has the following definition: consider a set of $M$ vertices and a parameter $\beta$ such that $10<\beta<$ 1, the probability of any two vertices being connected is $\beta$. This network has been used frequently as a first prototype to study the influence of the network topology on social dynamics [46, 47, 48].
The degree $k_{i}$ is the number of neighbors of node $i$ and the average degree $\mu$ of the network is defined as $\mu=(1/M)\sum_{i}^{M}k_{i}$ [49]. In the case of ER network, we have $\mu=\beta(M-1)$ [4]. So we use $\mu$ as control parameter of the network instead of $\beta$.
3 Results
@@
  • The rewiring happens with probability of $r^{q}$ where $L$ links are removed and new $L$ links are created involving the target agent.
@@
Figure 2: Results of the Reference Case on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ as function of $M$ for different average degrees $\mu$ of the ERN with no coevolution. These results is the reference to which we will compare the results with coevolution to evaluate its impact. Parameter: $q=1$ and $S=5\times 10^{4}$ samples. The green line is the IS (Eq. 4), black dashed line is the FCN result and the pink line is the approximation for large $M$. Purple, blue, green, orange and red circles correspond to an average degree of $\mu=1$, 5, 10, 20 and 100, respectively. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
Figure 3: Snaphots of instances of the random network used in this work. System size: $M=67$ agents. Each component is represented in a different color. (a) Low average degree scenario: $\mu=1.0$. The network is not percolated and the largest component is in painted in red. (b) High average degree scenario: $\mu=10.0$. In this scenario the network is percolated.
@@ 3.2 Reference case
Firstly we present the results with a static network: the Reference Case (RC). It is important to understand this case so we can highlight the effect of the coevolution. At $K=0$ (Figure 2(a)) the best result (lower $\left\langle C\right\rangle$) is given by the full connected network (FCN) and the worst by the independent search (IS). The increasing average degree $\mu$ interpolates from IS to the FCN case. As there is only one maximum (the global one) it becomes easier to increase the fitness with more neighbors to copy. However, in terms of system size $M$ there is one optimal value roughly between 15 and 20 which returns $\left\langle C\right\rangle\sim 0.007$. So, if there is too few neighbors (smaller $M$) or too much neighbors (larger $M$) the agents spend more time looking for the global maximum. At large $M\gtrsim 10^{4}$ the system behaviors as the limit case $M/2^{N}$.
Figure 4: Coevolution results in a low degree network ($\mu=1$) on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of modification $L$, limited by $L<k_{i}$ for each agent. $L=0$ means no coevolution. Parameters: $S=10^{5}$ samples and $\mu=1$. The green line is the IS (Eq. 4) and the black dashed line is the FCN result. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.

### v3

@@ Abstract
Coevolution on social models couples the time evolution of the network with the time evolution of the states of the agents. This paper presents a new coevolution dynamic allowing more than one rewiring on the network. We explore how this coevolution can be employed as an optimization strategy for problem-solving capability of task-forces. We use an agent-based model to study how this new coevolution dynamic can help a group of agents whose task is to find the global maxima of NK fitness landscapes. Each agent can replace more than one neighbor, and this quantity is a tunable parameter in the model. These rewirings is a way for the agent to obtain information from individuals that were not previously part of its neighborhood. Our results showed that this tunable coevolution can indeed produce gain on the computational cost under certain circunstances. At high average degree network and difficult landscape, the effect is complex. If the agent has a low fitness, 3 or 4 rewirings can bring some improvement.
@@ 1 Introduction
In society, it is common to face problems that require collaboration with other people, from everyday challenges to complex tasks, such as group projects at work. In this context, the search for more effective problem-solving strategies becomes a topic of great interest. Social interactions are inherently complex such that, in recent years, numerous models of social behavior have emerged to explore this intricacy [1, 2, 3] using the framework of Network Science [4]. The nodes of the network represent individuals (or agents) with states defined by the social model while the connections represent the interactions between different individuals. The states of the individuals evolve following the set of rules from the dynamic social model in study. The network itself can also have a time evolution: links can be destroyed or created.
@@
This work is organized as follows: In Section 2.1 we present a brief discussion about the NK model, in section 2.2 we define the computational cost, in section 2.3 we define how the agents are updated, in sections 2.4 and 2.5 we define the proposed coevolution mechanism and in section 2.6 we described the random network used to connect the agents.Section 3 details some required methods to analyze the results. Section 4 is dedicated to present and discuss the results and section 5 is dedicated to our final considerations.
@@ 2.1 NK Fitness Landscapes
The function $f(\mathcal{X})$ has a global maximum $f(\mathcal{X}_{G})=F$ at the string $\mathcal{X}_{G}$. The number of local maxima of the function $f(\mathcal{X})$ is zero at $K=0$ landscape and increases at $K>0$. Therefore, at $K=0$, the fitness difference $f(\mathcal{X}_{G})-f(\mathcal{X})$ is proportional to the Hamming distance $d_{H}(\mathcal{X}_{G},\mathcal{X})$ between the vectors [35]. The Hamming distance gives the number of different bits between the two vectors. For example, consider $\mathcal{X}_{1}=(010101)$ and $\mathcal{X}_{2}=(000101)$. The hamming distance is $d_{H}(\mathcal{X}_{1},\mathcal{X}_{2})=1$. At $K>0$, the landscape becomes more complicated because the function $f$ has additional local maxima beside the global one. Thus the difference $f(\mathcal{X}_{G})-f(\mathcal{X})$ is no longer proportional to the hamming distance $d_{H}$. The average number of local maxima increases with $K$ until the limiting case of $K=N-1$, in which the landscape becomes completely random, known as the Random Energy model [36].
@@ 2.2 Computational cost
The efficiency of the search is measured by the time $t_{g}$ in MCS required by the first agent who finds the global maximum $F$, in other words, when its string becomes $\mathcal{X}_{G}$ after a number of updates. A small $t_{g}$ means more efficient search: the goal was achieved in less time. So, the lower the better. The normalized computational cost is then defined as [11]:
@@
[…]
@@ 2.3 Strings update
Each agent has its string $\mathcal{X}$ updated when it is selected for analysis. This is how each agent tries to improve its fitness $f(\mathcal{X})$, to eventually reaches the global maximum $F$, which happens when $\mathcal{X}=\mathcal{X}_{G}$. Some basic definitions we will use throughout the text are:
  • Target: selected agent for the analysis, with string $\mathcal{X}_{t}$ and fitness $\phi_{t}=f(\mathcal{X}_{t})$.
  • Influence neighborhood: set of agents who are neighbors of the target $\mathcal{X}_{t}$.
  • Model $\mathcal{X}_{m}$: agent from this neighborhood with the highest fitness, with fitness $\phi_{m}=f(\mathcal{X}_{m})$.
  • Worst neighbor $\mathcal{X}_{s}$: agent on the neighborhood with the smallest fitness. We write its fitness as $\phi_{s}=f(\mathcal{X}_{s})$.
The target $\mathcal{X}_{t}$ may be the model $\mathcal{X}_{m}$ or the one with the smallest fitness, $\mathcal{X}_{s}$, so that we have $\phi_{s}\leq\phi_{t}\leq\phi_{m}$.
@@
The target can implement one of the two strategies to perform the update of its string $\mathcal{X}_{t}$: exploitation or exploration [33]. Exploitation implies the target tries to learn the solution from a neighbour with a higher fitness. This is implemented as a copy, with probability $p$: the target copies one bit of the model string $\mathcal{X}_{m}$ [35]. Otherwise, exploration means the target aims to innovate a new solution with a better fitness by itself. It is a mutation on the string of the target, and it should happen with probability $1-p$. However if the target has no neighbors or if it is the model of its influence neighborhood, $\mathcal{X}_{t}=\mathcal{X}_{m}$, the mutation will happen for sure. If it has no neighbors, it does not have anyone to copy. And if it is already the model, there is no point in copy another agent with a smaller (worst) fitness.
@@
Both processes change the string of the target agent. The mutation is achieved by a flip of a random bit of the string. The copy process involves more steps: the different bits bewteen the target and the model strings, $\mathcal{X}_{t}$ and $\mathcal{X}_{m}$, are identified. Then, one of them is randomly selected and flipped on the target string $\mathcal{X}_{t}$. This is the imperfect copy on the imitation learning context [31]. As a result, the Hamming distance $d_{H}(\mathcal{X}_{t},\mathcal{X}_{m})$ between the two strings decreases. This means the target agent increases his fitness, as $f(\mathcal{X}_{m})>f(\mathcal{X}_{t})$. If $K=0$, it means the target gets closer to the global maximum. On the other hand, if $K>0$, the model may be a local maxima and the target $\mathcal{X}_{t}$ can get further away from the global maximum $\mathcal{X}_{G}$: $d_{H}(\mathcal{X}_{t},\mathcal{X}_{G})$ increases while $d_{H}(\mathcal{X}_{t},\mathcal{X}_{m})$ decreases. The target can get trapped on the suboptimal solution (local maximum) [18]. $K=0$ is the easy landscape and increasing $K$ increases the difficulty of the landscape. The most difficult landscape is $K=N-1$, where there is no point on copy because the landscape is completely random.
Figure 1: Schematic illustration of the implemented model. (a) Illustration of one iteration. It starts with the selection of the target agent. $r$ is calculated through Eq. 6. With probability $r$ the network can suffer a rewiring. The next step is the string update, where a copy happens with probability $p$ and a mutation with probability $1-p$. (b) Illustration of the rewiring process. The target agent (indicated by a dashed black square) loses $L$ neighbors and gets new $L$ ones. The case $L=1$ is illustrated in this picture: the target loses the neighbor on the right and gets acquainted with the agent right above him. The colors of the nodes indicate the components.
Figure 2: Snaphots of instances of the random network used in this work. System size: $M=67$ agents. Each component is represented in a different color and the largest one is in painted in red. (a) Low average degree scenario: $\mu=1.0$. The network is not percolated. (b) High average degree scenario: $\mu=10.0$. In this scenario the network is percolated.
@@ 2.4 Coevolution mechanism
The rewiring itself is the temporal evolution of the network. To be considered as coevolution, this network evolution must be coupled to the evolution of the agent string. The constant rewiring (always happening) is not coevolution, although it is a network evolution [35]. As the rewiring itself is a random process, the coupling of the network evolution with the agent states, in our model, will be achieved by the probability of rewiring $r$: it must be calculated considering the fitness $\phi$ of the agents. The rewiring process happens before the string update, as illustrated on Fig. 1(a). Another illustration of one example of the rewiring itself (with $L=1$) is shown Fig. 1(b).
@@ 2.5 Rewiring probability
The goal of each agent is to increase its fitness $f(\mathcal{X})$ (given by Eq. 1) until one of them, eventually, reaches the global maximum fitness $F$ of the landscape. So, having neighbors with better fitness if beneficial. One can propose that the probability $r$ should be larger when most of the neighbors have smaller fitness, so one of them may be replaced with a higher fitness one. The limit cases of this hypothesis are:
  • the target is the model $\mathcal{X}_{t}=\mathcal{X}_{m}$: all neighbors have smaller fitness, so it is justifiable to replace some of them. The probability should be maximum: $r=1.0$.
  • the target is the one with the smallest fitness of the neighborhood $\mathcal{X}_{t}=\mathcal{X}_{s}$: any copy is already rewarding, so there is no urgency to replace any neighbor. The probability should be minimum: $r=0.0$.
  • Otherwise, the probability shoud be such as $1.00.0<r<$ 1.0.
@@
$r_{1}=\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}},$ (5)
where $\phi_{t}=f(\mathcal{X}_{t})$ is the fitness of the target agent $\mathcal{X}_{t}$, $\phi_{s}=f(\mathcal{X}_{s})$ is the smallest fitness and $\phi_{m}=f(\mathcal{X}_{m})$ is the maximum one (fitness of the model $\mathcal{X}_{m}$) on the influence neighborhood.
We have tested other definitions for the probability and, surprinsingly, the following one gave the most interesting results:
$r_{2}=1-\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}}.$ (6)
This is the opposite of the first attempt: $r=0.0$ if the target is the model and 1.0 if it has the smallest fitness. The two probabilities well be studied on the Results section.
@@ 2.6 Random Network
A network or a graph comprises a collection of vertices interconnected by a set of edges [38, 39]. Each edge, or connection, links a pair of vertices with the number of vertices defining the network size and the set of edges determining its connectivity. Vertices that share an edge are termed neighbors. Throughout this work the terms vertices, agents, and individuals have the same meaning. Various types of networks are used for solutions of physical and biological problems [40, 41, 42]. In this work we will use the Erdös-Rényi (ER) network [43, 44] which has the following definition: consider a set of $M$ vertices and a parameter $\beta$ such that $1.00.0<\beta<$ 1.0, the probability of any two vertices being connected is $\beta$. This network has been used frequently as a first prototype to study the influence of the network topology on social dynamics [45, 46, 47].
The degree $k_{i}$ is the number of neighbors of node $i$ and the average degree $\mu$ of the network is defined as $\mu=(1/M)\sum_{i}^{M}k_{i}$ [48]. In the case of ER network, we have $\mu=\beta(M-1)$ [4]. So we use $\mu$ as control parameter of the network instead of $\beta$. Figure 2 shows a graphical representation of the ER network with two different values of $\mu$.
3 Methods
@@
  • The rewiring happens with probability of $r$ where $L$ links are removed and new $L$ links are created involving the target agent.
@@
Figure 3: Results of the Reference Case on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ as function of $M$ for different average degrees $\mu$ of the ERN with no coevolution. These results is the reference to which we will compare the results with coevolution to evaluate its impact. Parameter: $S=5\times 10^{4}$ samples. The green line is the IS (Eq. refindsearch), black dashed line is the FCN result and the pink line is the approximation for large $M$. Purple, blue, green, orange and red circles correspond to an average degree of $\mu=1$, 5, 10, 20 and 100, respectively. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
@@ 3.2 Reference case
Firstly we present the results with a static network: the Reference Case (RC). It is important to understand this case so we can highlight the effect of the coevolution. At $K=0$ (Figure 3(a)) the best result (lower $\left\langle C\right\rangle$) is given by the full connected network (FCN) and the worst by the independent search (IS). The increasing average degree $\mu$ interpolates from IS to the FCN case. As there is only one maximum (the global one) it becomes easier to increase the fitness with more neighbors to copy. However, in terms of system size $M$ there is one optimal value roughly between 15 and 20 which returns $\left\langle C\right\rangle\sim 0.07$. So, if there is too few neighbors (smaller $M$) or too much neighbors (larger $M$) the agents spend more time looking for the global maximum. At large $M\gtrsim 10^{4}$ the system behaviors as the limit case $M/2^{N}$.
Figure 4: Coevolution results using $r_{1}$ (Eq. 5) as rewiring probability on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of modification $L$, limited by $L<k_{i}$ for each agent. $L=0$ means no coevolution. Parameters: $S=10^{5}$ samples and $\mu=10$. The green line is the IS (Eq. 4) and the black dashed line is the FCN result. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
Figure 5: Coevolution results using $r_{2}$ (Eq. 6) as rewiring probability on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of modification $L$, limited by $L<k_{i}$ for each agent. $L=0$ means no coevolution. Parameters: $S=10^{5}$ and $\mu=10$. The green line is the IS (Eq. 4) and the black dashed line is the FCN result. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
@@
Figure 6: 2D map of the computational cost $\left\langle C\right\rangle$ and its gain $\Delta$ in the $(L,\mu)$ space. Parameters: $K=4$, $M=67$ and $S=10^{5}$. (a) $\log\left\langle C\right\rangle$. (b) $\Delta$ defined on Eq. 8. Positive values are indicated in red and negative values in blue.
4 Results